OpenStudy (anonymous):

Whoever sets this up and shows the correct steps and answer gets a medal and fan. Ray SP bisects

OpenStudy (anonymous):

SP bisects <RST m<RST= m<PST 5/2y = (y+5) y= 10/3 = 3.333333333

OpenStudy (anonymous):

how'd you get 10/3 again?